# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ B.1 Linearization
$I_{2}\leq\frac{1}{n}(B_{x}^{2}+1)\|\boldsymbol{a}(t)\|_{2}^{2}\kappa_{2}(C_{1})=O_{p}(n^{-1})\kappa_{2}(C_{1}),\quad I_{3}\leq O_{p}(n^{-1})\kappa_{2}(C_{1}).$
@@
Recall $\kappa_{1}(C_{1})=\inf_{z\in[-C_{1},C_{1}]}(\phi^{\prime\prime}(z))^{-1}$. Let $\mathbf{r}(t)=\mathbf{f}(t)-\tilde{\mathbf{f}}(t)$ and $C_{2}=(\kappa_{1}(C_{1})\lambda_{0})/4$. Consider the function $u(t)=\|\mathbf{r}(t)\|_{2}$ defined on $t\in[0,t_{1}]$. Clearly, $u(t)$ is continuous on $[0,t_{1}]$. Since $u(t)=0$ if and only if $\mathbf{r}(t)=0$, we have that $u(t)$ is differentiable whenever $u(t)\neq 0$. Especially, when $u(t)\neq 0$, notice that
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))^{2}$ $=\frac{\mathrm{d}}{\mathrm{dt}}\|e^{C_{2}\eta_{0}t}\mathbf{r}(t)\|_{2}^{2}$ (84)
$=2e^{C_{2}\eta_{0}t}\mathbf{r}(t)^{T}\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}\mathbf{r}(t))$
$=2e^{2C_{2}\eta_{0}t}\mathbf{r}(t)^{T}\Big{(}C_{2}\eta_{0}\mathbf{r}(t)-\eta_{0}H(t)(\mathbf{f}(t)-\mathbf{y})+\eta_{0}\tilde{H}(t)(\tilde{\mathbf{f}}(t)-\mathbf{y})\Big{)}$
$=2\eta_{0}e^{2C_{2}\eta_{0}t}\Big{(}\mathbf{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\mathbf{f}}(t)-\mathbf{y})-\mathbf{r}(t)^{T}(H(t)-C_{2}I_{n})\mathbf{r}(t)\Big{)}$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}\mathbf{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\mathbf{f}}(t)-\mathbf{y})$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}u(t)\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2},$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))\leq\eta_{0}e^{C_{2}\eta_{0}t}\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2}.$
Now consider $t\in[0,t_{1}]$. Let $t^{\prime}=\inf\{s\in[0,t]\colon\mathbf{r}(s)\neq 0\}$. Then we have
$e^{C_{2}\eta_{0}t}u(t)\leq\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s.$
@@
$\|\mathbf{r}(t)\|_{2}$ $\leq e^{-C_{2}\eta_{0}t}\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s$ (85)
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}(5mK^{2}(B_{x}+1)^{2}LCn^{-1/2})\int_{t^{\prime}}^{t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}\mathrm{d}s$
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}5mK^{2}(B_{x}+1)^{2}LC\|\mathbf{f}(0)-\mathbf{y}\|_{2}tn^{-1/2}.$
@@
$=$ $\frac{\eta_{0}}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(na_{k}(t))}J_{a_{k}}\mathbf{f}(t)^{T}(\mathbf{f}(t)-\mathbf{y})-\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}J_{a_{k}}\mathbf{f}(0)^{T}(\tilde{\mathbf{f}}(t)-\mathbf{y})\Big{|}$
$\leq$ $\frac{\eta_{0}}{n}\Big{(}\left|\frac{1}{\phi^{\prime\prime}(na_{k}(t))}-\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\mathbf{f}(t)\|_{2}\|\mathbf{f}(t)-\mathbf{y}\|_{2}$
$\quad+\left|\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\mathbf{f}(t)-J_{a_{k}}\mathbf{f}(0)\|_{2}\|\mathbf{f}(t)-\mathbf{y}\|_{2}$
$\quad+\left|\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\mathbf{f}(0)\|_{F}\|\mathbf{r}(t)\|_{2}\Big{)}$
@@
$\leq$ $\frac{\eta_{0}}{n}\Big{(}LCn^{-1/2}\cdot\sqrt{m}(B_{x}|w_{k}(t)|+|b_{k}(t)|)\cdot\|\mathbf{f}(0)-\mathbf{y}\|_{2}e^{-C_{2}\eta_{0}t}$
$+\kappa_{2}(C_{1})\cdot\sqrt{m}(B_{x}|w_{k}(t)-w_{k}(0)|+|b_{k}(t)-b_{k}(0)|)\cdot\|\mathbf{f}(0)-\mathbf{y}\|_{2}e^{-C_{2}\eta_{0}t})$
$+\kappa_{2}(C_{1})\cdot\sqrt{m}(B_{x}|w_{k}(0)|+|b_{k}(0)|)\cdot\eta_{0}e^{-C_{2}\eta_{0}t}5mK^{2}(B_{x}+1)^{2}LC\|\mathbf{f}(0)-\mathbf{y}\|_{2}tn^{-1/2}\Big{)}$
$\leq$ $\frac{2\eta_{0}\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\exp(C_{2}\eta_{0}t)}\Big{(}LC\cdot K(B_{x}+1)n^{-3/2}+O_{p}(n^{-2})\kappa_{2}(C_{1})C$
@@
$\leq$ $\frac{2LC(B_{x}+1)K\eta_{0}\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\exp(C_{2}\eta_{0}t)}\big{(}2+5\kappa_{2}(C_{1})(B_{x}+1)^{2}K^{2}\eta_{0}mt\Big{)}n^{-3/2}$
@@
$<$ $2LC(B_{x}+1)K\eta_{0}\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}\Big{(}\int_{0}^{\infty}\frac{2+5\kappa_{2}(C_{1})(B_{x}+1)^{2}K^{2}\eta_{0}ms}{\exp(C_{2}\eta_{0}s)}\mathrm{d}s\Big{)}n^{-3/2}$
$=$ $2LC(B_{x}+1)K\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}\frac{8\kappa_{1}(C_{1})\lambda_{0}+80\kappa_{2}(C_{1})m(B_{x}+1)^{2}K^{2}}{(\kappa_{1}(C_{1})\lambda_{0})^{2}}n^{-3/2}$
$\leq$ $\kappa_{2}(C_{1})^{2}\kappa_{3}(C_{1})\frac{\kappa_{1}(C_{1})\lambda_{0}+10m(B_{x}+1)^{2}K^{2}\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})^{2}}\cdot\frac{16(B_{x}+1)K\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{(\lambda_{0})^{2}}Cn^{-3/2}$
@@
$\|\mathbf{f}(t)-\tilde{\mathbf{f}}(t)\|_{2}=o_{p}(1)e^{-C_{2}\eta_{0}t}t.$ (86)
Next we fix the constants $C_{1}$ and $C$ and aim to show $f(\cdot,\theta(t))$ agrees with $f(\cdot,\tilde{\theta}(t))$ on any given test data $x\in\mathbb{R}$. For a fix $x\in\mathbb{R}$, consider
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\theta(t))$ $=-\eta_{0}H(t,x)(\mathbf{f}(t)-\mathbf{y})$
$H(t,x)$ $\triangleq\nabla_{\theta}f(x,\theta(t))(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}\mathbf{f}(t)^{T}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\tilde{\theta}(t))$ $=-\eta_{0}\tilde{H}(t,x)(\tilde{\mathbf{f}}(t)-\mathbf{y})$
$\tilde{H}(t,x)$ $=\nabla_{\boldsymbol{a}}f(x,\hat{\theta})(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\boldsymbol{a}}\mathbf{f}(0)^{T}$
@@
Notice that for large enough $n$ we have
$\|J_{\theta}\mathbf{f}(t)\|_{F}^{2}$ $=\sum_{i=1}^{m}\Big{(}1+\sum_{k=1}^{n}|\sigma(w_{k}(t)x_{i}-b_{k}(t))|^{2}+\sum_{k=1}^{n}|a_{k}(t)|^{2}|\sigma^{\prime}(w_{k}(t)x_{i}-b_{k}(t))|^{2}(x_{i}^{2}+1)\Big{)}$ (88)
$\leq m+\sum_{i=1}^{m}\sum_{k=1}^{n}\Big{(}2|w_{k}(t)|^{2}|x_{i}|^{2}+2|b_{k}(t)|^{2}+|a_{k}(t)|^{2}(|x_{i}|^{2}+1)\Big{)}$
$\leq m(1+2B_{x}^{2}\|\boldsymbol{w}(t)\|_{2}^{2}+2\|\boldsymbol{b}(t)\|_{2}^{2}+(B_{x}^{2}+1)\|\boldsymbol{a}(t)\|_{2}^{2}).$
By (82) and $\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1}$ we have $\|J_{\theta}\mathbf{f}(t)\|_{2}=O_{p}(\sqrt{n})$. Note a similar bound applies to $\|J_{\theta}f(x,\theta(t))\|_{2}$. Then with large enough $n$ we have for $t\geq 0$,
$\|H(t,x)\|_{2}\leq\frac{1}{n}\|J_{\theta}f(x,\theta(t))\|_{2}\|Q(t)\|_{2}\|J_{\theta}\mathbf{f}(t)\|_{2}=O_{p}(1).$
@@
$\leq\int_{0}^{t}\Big{|}\eta_{0}H(s,x)(\mathbf{f}(s)-\mathbf{y})-\eta_{0}\tilde{H}(s,x)(\tilde{\mathbf{f}}(s)-\mathbf{y})\Big{|}\mathrm{d}s$
$\leq\int_{0}^{t}\eta_{0}\|H(s,x)\|_{2}\|\mathbf{f}(s)-\tilde{\mathbf{f}}(s)\|_{2}$
$\quad+\eta_{0}\|H(s,x)-\tilde{H}(s,x)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s$
$\leq\int_{0}^{\infty}O_{p}(1)o_{p}(1)e^{-C_{2}\eta_{0}s}s\mathrm{d}s+O_{p}(n^{-1/2})\int_{0}^{\infty}\|\mathbf{f}(0)-\mathbf{y}\|_{2}^{2}e^{-C_{2}\eta_{0}s}\mathrm{d}s$
@@
Next we show that, if the potential $\Phi$ satisfies Assumption 7, one can select a sufficiently large $\omega$ to ensure that the conditions in Proposition 24 and 26 are satisfied.
@@
$D_{1}=\frac{8\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}},\quad D_{2}=10m(B_{x}+1)^{2}K^{2}.$
@@
In the following result, we show that when $\phi$ is the hypentropy potential $\rho_{\beta}$, the conditions in Proposition 24 and 26 are satisfied with sufficiently large $\beta$.
@@
Therefore, by selecting $C_{1}=2\frac{8\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}}$, we have that for sufficiently large $\beta$,
$\frac{C_{1}\kappa_{1}(C_{1})}{\kappa_{2}(C_{1})}\geq\frac{C_{1}}{2}=\frac{8\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}}.$

### v2

@@ B.1 Linearization
$I_{2}\leq\frac{d}{n}(B_{x}^{2}+1)\|\boldsymbol{a}(t)\|_{2}^{2}\kappa_{2}(C_{1})=O_{p}(n^{-1})\kappa_{2}(C_{1}),\quad I_{3}\leq O_{p}(n^{-1})\kappa_{2}(C_{1}).$
@@
Recall $\kappa_{1}(C_{1})=\inf_{z\in[-C_{1},C_{1}]}(\phi^{\prime\prime}(z))^{-1}$. Let $\boldsymbol{r}(t)=\boldsymbol{f}(t)-\tilde{\boldsymbol{f}}(t)$ and $C_{2}=(\kappa_{1}(C_{1})\lambda_{0})/4$. Consider the function $u(t)=\|\boldsymbol{r}(t)\|_{2}$ defined on $t\in[0,t_{1}]$. Clearly, $u(t)$ is continuous on $[0,t_{1}]$. Since $u(t)=0$ if and only if $\boldsymbol{r}(t)=0$, we have that $u(t)$ is differentiable whenever $u(t)\neq 0$. In particular, when $u(t)\neq 0$, notice that
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))^{2}$ $=\frac{\mathrm{d}}{\mathrm{dt}}\|e^{C_{2}\eta_{0}t}\boldsymbol{r}(t)\|_{2}^{2}$ (87)
$=2e^{C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}\boldsymbol{r}(t))$
$=2e^{2C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}\Big{(}C_{2}\eta_{0}\boldsymbol{r}(t)-\eta_{0}H(t)(\boldsymbol{f}(t)-\boldsymbol{y})+\eta_{0}\tilde{H}(t)(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})\Big{)}$
$=2\eta_{0}e^{2C_{2}\eta_{0}t}\Big{(}\boldsymbol{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})-\boldsymbol{r}(t)^{T}(H(t)-C_{2}I_{n})\boldsymbol{r}(t)\Big{)}$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}u(t)\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2},$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))\leq\eta_{0}e^{C_{2}\eta_{0}t}\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2}.$
Now consider $t\in[0,t_{1}]$. Let $t^{\prime}=\inf\{s\in[0,t]\colon\boldsymbol{r}(s)\neq 0\}$. Then we have
$e^{C_{2}\eta_{0}t}u(t)\leq\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s.$
@@
$\|\boldsymbol{r}(t)\|_{2}$ $\leq e^{-C_{2}\eta_{0}t}\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s$ (88)
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}(5mK^{2}(B_{x}+1)^{2}LCn^{-1/2})\int_{t^{\prime}}^{t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}\mathrm{d}s$
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}5mK^{2}(B_{x}+1)^{2}LC\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}tn^{-1/2}.$
@@
$=$ $\frac{\eta_{0}}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(na_{k}(t))}J_{a_{k}}\boldsymbol{f}(t)^{T}(\boldsymbol{f}(t)-\boldsymbol{y})-\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}J_{a_{k}}\boldsymbol{f}(0)^{T}(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})\Big{|}$
$\leq$ $\frac{\eta_{0}}{n}\Big{(}\left|\frac{1}{\phi^{\prime\prime}(na_{k}(t))}-\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\boldsymbol{f}(t)\|_{2}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}$
$\quad+\left|\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\boldsymbol{f}(t)-J_{a_{k}}\boldsymbol{f}(0)\|_{2}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}$
$\quad+\left|\frac{1}{\phi^{\prime\prime}(n\tilde{a}_{k}(t))}\right|\|J_{a_{k}}\boldsymbol{f}(0)\|_{F}\|\boldsymbol{r}(t)\|_{2}\Big{)}$
@@
$\leq$ $\frac{\eta_{0}}{n}\Big{(}LCn^{-1/2}\cdot\sqrt{m}(B_{x}\|w_{k}(t)\|_{2}+|b_{k}(t)|)\cdot\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}e^{-C_{2}\eta_{0}t}$
$+\kappa_{2}(C_{1})\cdot\sqrt{m}(B_{x}\|w_{k}(t)-w_{k}(0)\|_{2}+|b_{k}(t)-b_{k}(0)|)\cdot\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}e^{-C_{2}\eta_{0}t})$
$+\kappa_{2}(C_{1})\cdot\sqrt{m}(B_{x}\|w_{k}(0)\|_{2}+|b_{k}(0)|)\cdot\eta_{0}e^{-C_{2}\eta_{0}t}5mK^{2}(B_{x}+1)^{2}LC\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}tn^{-1/2}\Big{)}$
$\leq$ $\frac{2\eta_{0}\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\exp(C_{2}\eta_{0}t)}\Big{(}LC\cdot K(B_{x}+1)n^{-3/2}+O_{p}(n^{-2})\kappa_{2}(C_{1})C_{1}$
@@
$\leq$ $\frac{2LC(B_{x}+1)K\eta_{0}\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\exp(C_{2}\eta_{0}t)}\big{(}2+5\kappa_{2}(C_{1})(B_{x}+1)^{2}K^{2}\eta_{0}mt\Big{)}n^{-3/2}$
@@
$<$ $2LC(B_{x}+1)K\eta_{0}\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}\Big{(}\int_{0}^{\infty}\frac{2+5\kappa_{2}(C_{1})(B_{x}+1)^{2}K^{2}\eta_{0}ms}{\exp(C_{2}\eta_{0}s)}\mathrm{d}s\Big{)}n^{-3/2}$
$=$ $2LC(B_{x}+1)K\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}\frac{8\kappa_{1}(C_{1})\lambda_{0}+80\kappa_{2}(C_{1})m(B_{x}+1)^{2}K^{2}}{(\kappa_{1}(C_{1})\lambda_{0})^{2}}n^{-3/2}$
$\leq$ $\kappa_{2}(C_{1})^{2}\kappa_{3}(C_{1})\frac{\kappa_{1}(C_{1})\lambda_{0}+10m(B_{x}+1)^{2}K^{2}\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})^{2}}\cdot\frac{16(B_{x}+1)K\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{(\lambda_{0})^{2}}Cn^{-3/2}$
@@
$\|\boldsymbol{f}(t)-\tilde{\boldsymbol{f}}(t)\|_{2}=o_{p}(1)e^{-C_{2}\eta_{0}t}t.$ (89)
Next we fix the constants $C_{1}$ and $C$ and aim to show $f(\cdot,\theta(t))$ agrees with $f(\cdot,\tilde{\theta}(t))$ on any given test data $x\in\mathbb{R}^{d}$. For a fix $x\in\mathbb{R}^{d}$, consider
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\theta(t))$ $=-\eta_{0}H(t,x)(\boldsymbol{f}(t)-\boldsymbol{y})$
$H(t,x)$ $\triangleq\nabla_{\theta}f(x,\theta(t))(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}\boldsymbol{f}(t)^{T}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\tilde{\theta}(t))$ $=-\eta_{0}\tilde{H}(t,x)(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})$
$\tilde{H}(t,x)$ $=\nabla_{\boldsymbol{a}}f(x,\hat{\theta})(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\boldsymbol{a}}\boldsymbol{f}(0)^{T}$
@@
For large enough $n$, by (85) and $\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1}$, we have
$\|J_{\theta}\boldsymbol{f}(t)\|_{F}^{2}$ $=\sum_{i=1}^{m}\Big{(}1+\sum_{k=1}^{n}|\sigma_{ik}(t)|^{2}+\sum_{k=1}^{n}|a_{k}(t)|^{2}|\sigma_{ik}^{\prime}(t)|^{2}(\|x_{i}\|_{2}^{2}+1)\Big{)}$ (91)
$\leq m+\sum_{i=1}^{m}\sum_{k=1}^{n}\Big{(}2\|w_{k}(t)\|_{2}^{2}\|x_{i}\|^{2}+2|b_{k}(t)|^{2}+|a_{k}(t)|^{2}(\|x_{i}\|^{2}+1)\Big{)}$
$=O_{p}(n),$
and hence $\|J_{\theta}\boldsymbol{f}(t)\|_{2}=O_{p}(\sqrt{n})$. Note a similar bound applies to $\|J_{\theta}f(x,\theta(t))\|_{2}$. Then with large enough $n$ we have for $t\geq 0$,
$\|H(t,x)\|_{2}\leq\frac{1}{n}\|J_{\theta}f(x,\theta(t))\|_{2}\|Q(t)\|_{2}\|J_{\theta}\boldsymbol{f}(t)\|_{2}=O_{p}(1).$
@@
$\leq\int_{0}^{t}\Big{|}\eta_{0}H(s,x)(\boldsymbol{f}(s)-\boldsymbol{y})-\eta_{0}\tilde{H}(s,x)(\tilde{\boldsymbol{f}}(s)-\boldsymbol{y})\Big{|}\mathrm{d}s$
$\leq\int_{0}^{t}\eta_{0}\|H(s,x)\|_{2}\|\boldsymbol{f}(s)-\tilde{\boldsymbol{f}}(s)\|_{2}$
$\quad+\eta_{0}\|H(s,x)-\tilde{H}(s,x)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s$
$\leq\int_{0}^{\infty}O_{p}(1)o_{p}(1)e^{-C_{2}\eta_{0}s}s\mathrm{d}s+O_{p}(n^{-1/2})\int_{0}^{\infty}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}^{2}e^{-C_{2}\eta_{0}s}\mathrm{d}s$
@@
Next we show that, if the potential $\Phi$ satisfies Assumption 8, one can select a sufficiently large $\omega$ to ensure that the conditions in Proposition 27 and 29 are satisfied.
@@
$D_{1}=\frac{8\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}},\quad D_{2}=10m(B_{x}+1)^{2}K^{2}.$
@@
In the following result, we show that when $\phi$ is the hypentropy potential $\rho_{\beta}$, the conditions in Proposition 27 and 29 are satisfied with sufficiently large $\beta$.
@@
Therefore, by selecting $C_{1}=2\frac{8\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}}$, we have that for sufficiently large $\beta$,
$\frac{C_{1}\kappa_{1}(C_{1})}{\kappa_{2}(C_{1})}\geq\frac{C_{1}}{2}=\frac{8\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}}.$